OpenStudy (anonymous):

HEEELPPP!!!!Solve by elimination. 1.7x + 3.5y = 0.01 0.6x + 1.2y = 0

OpenStudy (anonymous):

Select either eq and solve for y. I selected 0.6x +1.2y = 0 then 1.2y = -0.6x Divide both sides by 1.2 so y = -(0.6/1.2)x or y = -0.5x

OpenStudy (anonymous):

Use y = -0.5x to substitute for y in first equation 1.7x + 3.5(-0.5y) = 0.01 1.7y - 1.75y = 0.01 -0.05y = 0.01

OpenStudy (anonymous):

Are you still with me? Can you bring it home from here?

OpenStudy (anonymous):

yes, thank you so much!

OpenStudy (anonymous):

it says in my textbook that the answers are (-0.2, 0.1) .....

OpenStudy (anonymous):

Ouch, did you see my typo? Sorry about that. Should be 1.7x - 1.75x = 0.01 then -0.05x = 0.01 so x = -0.2 and y = 0.1

OpenStudy (anonymous):

Yes, these answers agree with your textbook

OpenStudy (anonymous):

Thanks for the medal and apologies for the typo.

OpenStudy (anonymous):

Thank you! I get it now :)

OpenStudy (anonymous):

It's alright, at least you helped.